I am trying to solve for x on this problem, except I keep getting stuck and can not seem to get this correct. the problem is: 2 - (5/3)x = 3x + (3/2), please let me know how you arrive at a solution if you would.
Thank you
Dave

Answer by Alan3354(69443)

2 - (5/3)x = 3x + (3/2)
Multiply by 6 to eliminate fractions
12 - 10x = 18x + 9
3 = 28x
x = 3/28
-----------
To check:
2 - (5/3)*(3/28) = 3*(3/28) + 3/2
56/28 - 5/28 = 9/28 + 42/28
47/28 = 47/28